Find the general solution of the equation of the equation cos3x_sinx=cos5x
 

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
Cos 3x - sin x = cos 5x
 
Cos 3x - cos 5x -  sinx = 0
 
2 sin 4x sin x - sinx = 0
 
Sinx = 0.  Or 2 sin4x - 1 = 0
 
x = n*pi.    Or 4x = n*pi + (-1)ⁿ pi/6
 
x = npi/4 + (-1)ⁿ pi/24